### v1

# Solving Distributed Flexible Job Shop Scheduling Problems in the Wool Textile Industry with Quantum Annealing

Abstract
Many modern manufacturing companies have evolved from a single production site to a multi-factory production environment that must handle both geographically dispersed production orders and their multi-site production steps. The availability of a range of machines in different locations capable of performing the same operation and shipping times between factories have transformed planning systems from the classic Job Shop Scheduling Problem (JSSP) to Distributed Flexible Job Shop Scheduling Problem (DFJSP). As a result, the complexity of production planning has increased significantly. In our work, we use Quantum Annealing (QA) to solve the DFJSP. In addition to the assignment of production orders to production sites, the assignment of production steps to production sites also takes place. This requirement is based on a real use case of a wool textile manufacturer. To investigate the applicability of this method to large problem instances, problems ranging from 50 variables up to 250 variables, the largest problem that could be embedded into a D-Wave quantum annealer Quantum Processing Unit (QPU), are formulated and solved. Special attention is dedicated to the determination of the Lagrange parameters of the Quadratic Unconstrained Binary Optimization (QUBO) model and the QPU configuration parameters, as these factors can significantly impact solution quality. The obtained solutions are compared to solutions obtained by Simulated Annealing (SA), both in terms of solution quality and calculation time. The results demonstrate that QA has the potential to solve large problem instances specific to the industry.
Keywords Quantum annealing Simulated annealing DFJSP Production planning
1 Introduction
This paper is based on a use case of a textile manufacturing company that operates in the worsted yarn industry and specializes in pure wool and wool blends for weaving, flat and circular knitting, luxury, as well as the non-apparel market. This business-to-business company has production sites in several countries on different continents. Production sites in different locations perform only specific production steps and work together in sequence-dependent setups to fulfill each individual sales order. For this purpose, sales orders first become production orders that must be assigned to the various production sites for the processing of various production steps. Given the scattering of production sites that might be involved in the processing of one single production order, a thorough production planning is critical.
The reason is that this production planning problem is an NP-hard combinatorial optimization problem, which can be reduced to the Distributed Flexible Job Shop Scheduling Problem (DFJSP) in this use case. It is widely known as one of the most challenging combinatorial optimization problems [15]. At the same time, operational production planning requires continuous and process-parallel optimization of machine allocation and processing sequence.
Quantum annealing (QA) is a promising technology to solve such difficult combinatorial optimization problems. While gate-based quantum computers can handle a broader range of problems, quantum annealers are specifically designed to solve combinatorial optimization problems commonly encountered in industry [26]. In this work, we use the AdvantageSystem4.1 quantum annealer from D-Wave to solve the DFJSP. To the best of our knowledge, there is no other work that specifically addresses solving the DFJSP using QA. Our contributions are as follows:
1.
We use a QA to solve the DFJSP. Due to the specifics of the wool textile industry, we also extend the DFJSP and make it even more complex: The classic DFJSP is extended to not only consider the distribution of production orders, but also the distribution of production steps (of a production order). Shipping times between production steps at different sites must also be included in the model. To the best of our knowledge, there is no other work in literature that discusses the solution of the classical or the extended DFJSP with QA.
2.
Calculation of Lagrange parameters and QPU configuration parameters. The values of the so-called Lagrange parameters of the QUBO model and the QPU configuration parameters have a significant influence on the solution quality. Instead of the trial-and-error method, as some approaches suggest111https://www.dwavesys.com/media/qvbjrzgg/guide-2.pdf, we determine these parameters based on the mathematical formulation of the problem.
3.
Assessment from an economic perspective. We address the question whether the use of QA can lead to a speed advantage for the DFJSP with distributed operations specific to the wool textile industry.
This paper is organized as follows: In Section 2 we review the state of the art on optimization problems and quantum annealing before we discuss related work in Section 3. We describe our QUBO model to solve the extended DFJSP in Section 4. Based on the QUBO model, we describe and discuss our experiments in Section 5. Section 6 provides a summary.
2 Preliminaries
In Subsection 2.1, we first study computationally complex combinatorial optimization problems that are typical in production planning. We then turn to the topic of quantum annealing in Subsection 2.2 and to simulated annealing in Subsection 2.3. The connection between Binary Quadratic Model (BQM) / Quadratic Unconstrained Binary Optimization (QUBO) and QA is highlighted in Subsection 2.4.
2.1 Combinatorial Optimization Problems
The Job Shop Scheduling Problem (JSSP) is a combinatorial optimization problem222Combinatorial optimization is the search process for maxima (or minima) of an objective function where the domain has a discrete but large configuration space. where a given set of jobs (or production orders) must be executed while only a limited number of machines can perform the operations of the jobs (also called production steps). Each operation can be executed by a specific machine. A schedule must be found that determines which jobs are to be executed at which time and by which machine. The aim is to minimize the total amount of time until all jobs have been completed [4, 1]. Many variations of the JSSP problem exist. To be closer, both to the market and manufacturing resources, modern enterprises have evolved from a single factory to a multi-factory manufacturing environment. This has led to the scheduling system changing from a JSSP system to a Flexible Job Shop Scheduling Problem (FJSP) and further to a Distributed Flexible Job Shop Scheduling Problem (DFJSP).
The FJSP is a generalization of the basic JSSP that allows to process tasks on one machine, chosen from a set of alternative machines (the machines in each set are identical) [19]. The FJSP is considered one of the most difficult combinatorial optimization problems, and it was revealed that it is strongly NP-hard since the classical JSSP has also proven to be strongly NP-hard [28]. A distinctive feature of JSSP (and FJSP) is the parallel job processing. While different process chains of an order are valuated, the choice of a process chain for an order affects the choice of production routes for all other orders.
The DFJSP refers to production activities that are carried out by several geographically distributed factories. The DFJSP can be regarded as an extension of FJSP which also includes the selection of suitable factories for the jobs [15]. The DFJSP described in the literature implies that if a job is assigned to a factory, then all the operations of this job must be processed at the same factory [13, 15]. The wool textile industry is a special case of DFJSP, where operation of a job can be processed in different locations. Assigning specific operations to different factories results in unique production schedules for the jobs, which influences the production planning. The DFJSP is more complex than the standard FJSP and is also classified as NP-Hard in complexity theory [15]. The goal of scheduling to assign each operation to factories and machines, considers the shipping time between locations, and launches operations on machines in a way that minimizes the makespan of all operations, thus ensuring that all jobs are completed in the shortest time possible.
2.2 Quantum Annealing
Today, two main architectures for quantum computers exist: gate model (also known as circuit model) and QA [4]. Gate model quantum computers use quantum gates to perform calculations, analogous to Boolean gates in classical computers. These processors are not yet fault-tolerant or large enough to reach quantum supremacy, a state that is referred to as the Noisy Intermediate-Scale Quantum (NISQ) era [3]. Consequently, it is difficult to develop practical applications on universal quantum gate model computers at present.
On the other hand, QA devices, like the computer developed by D-Wave, map a problem to a mathematical model of interacting magnets, then search for an optimal solution for problems with a large solution space. In general, QA is a mechanical metaheuristic to solve problems with a large, but discrete, search space.
QA systems are initialized in a low-energy state. So, the parameters of a problem are changed slowly, which allows the system to end in a low-energy state of the problem that corresponds to an optimal solution [5]. The evolution of the system can be represented by a time-dependent Hamiltonian333A classical Hamiltonian is a mathematical description of some physical system in terms of its energies [5].:
where A and B are time-dependent coefficients, s is the annealing time rescaled to take values between 0 and 1, are the Pauli Matrices444In quantum mechanics, Pauli Matrices are a set of 2 × 2 complex Hermitian and unitary matrices that arise in Pauli’s treatment of spins. of operations on a qubit q, h is the qubit bias and J the coupling strength between qubits. The algorithm gradually transitions in time from the initial ground state of the Initial Hamiltonian to a state described by the Final Hamiltonian [21]. The Final Hamiltonian matches the energy function to be minimized, and represents the best solution at its minimum value.
2.3 Simulated Annealing
Simulated Annealing (SA) is a nature inspired stochastic metaheuristic, single solution-based algorithm used to find a global optimum in a large problem space [22], and can be seen as a classical computing pendant to QA. Fig. 1 shows a comparison between QA and SA.
In SA, to simulate the annealing process, a temperature variable is set high and then the system is allowed to cool while the algorithm runs. SA is using thermal effects to converge to a low energy state, but therefore needs to jump over higher energy barriers. On the other hand, QA uses the quantum tunneling effect555In quantum mechanics, the tunneling effect describes the penetration of particles through barriers, even if particle energy is less than the barrier height [8]., which allows it to potentially surpass this higher barrier directly and thus reach the minimum faster. Recent advancements in QA methods have suggested potential benefits by employing QA over SA to address NP-hard optimization challenges [16].
2.4 Binary Quadratic Models
For QA, cost functions are mathematical expressions of the problem formulated as quadratic models that have the lowest values for good solutions. To solve problems on D-Wave quantum annealer, the problem space has to be formulated as a BQM, the D-Wave Systems software representation of the objective function:
where s and x are binary variables, i.e. categorical variables that can only take one of two values. s is the decision variable called spin that can either take values from the set {-1,1}, in which case the problem is called an Ising problem (Eq. 2) [14]. In the Ising model, h is the magnetic field that represents the linear coefficients corresponding to qubit biases and J represents the coupling strength between qubits [27]. In a QUBO problem (Eq. 3), x can take values from the set {0,1} [12]. Qubit biases are represented by the linear coefficients a and the coupling strength between qubits by the quadratic coefficients b. Both versions are NP-hard and can easily be converted into each other by applying variable substitution [25]. In this paper, the QUBO format is used to formulate the problem.
To solve a problem on a D-wave quantum annealer, the problem is formulated as an objective function in QUBO format. Low energy states of the objective function represent good solutions of the problem. Because the cost function can be represented as a graph, it can be mapped onto the QPU: linear coefficients to qubit biases and quadratic coefficients to coupler strengths. The QPU uses quantum annealing to seek the minimum in the resulting energy landscape, which corresponds to the solution of the problem. QA devices are less affected by noise than gate model quantum computers. This allows more qubit usage, and thus solves larger problems. QA provides an approach that promises good solutions to NP-hard problems and is specialized in combinatorial optimization. While the discussion continues whether QA can provide quantum speed-up over classical approaches, there have been several attempts to solve application-motivated problems with QA [26].
3 Related Work
In this Section, we assess previous approaches to solve the optimization problems described in Subsection 2.1.
JSSP. In the past a variety of approaches to solve the JSSP or its variation existed. Denkena et al. [9] mention heuristics, Genetic Algorithm (GA), Tabu Search, Ant Colony, Particle Swarm Optimization and SA. These approaches are used both individually and in hybrid combinations with one another. Venturelli et al. [24] present a solver for the JSSP that makes use of a QA. Carugno et al. [4] study the application of QA to solve JSSP and discuss challenges related to problem formulation to the fine-tuning of the QA. Carugno et al. [4] study the application of QA to solve the JSSP. In terms of solution quality, QA remarkably effective in comparison to classical approaches on very constrained problem instances, while closely competitive on realistic problem instances. Pakhomchik et al. [20] focus on the mathematical formulation of JSSP and propose a new large neighborhood search approach, with good empirical results for large problem instances. Zielewski et al. [29] use QA to solve JSSP and especially investigate the embedding process in the performance. The authors show that the initial setup of the problem, post-processing and problem formulation, as well as excess variables and large embeddings have a significant impact on results. A. Sharma and P. Maharjan [2] compare QA to SA through existing published works and conclude that QA has the potential to be considered as an optimized technique against SA. QA is more efficient than SA in estimating the ground-state of the system for small and for large sample size simulations. Denchev et al. [8] show that, for problems with tall and narrow energy barriers separating local minima, QA is more than times faster than SA running on a single core.
Overall, research considering transportation time focusses on flow shop scheduling problem666Flow shop scheduling is a special case of job shop scheduling in which all jobs have identical operation flow patterns on machines [1]. or classical FJSP. Most of the QA research in this field focuses on solving JSSP or FJSP and transportation time is often neglected in literature. However, transportation time between locations can have a big impact on maximum completion time of FJSP.
FSP. Flow Shop Scheduling Problem (FSP) is one of the most studied problems in order planning. Hurink and Knust [10] apply SA to minimize both total completion time and total tardiness of flow shop problems. Naderi et al. [17] propose an adaptation of the SA algorithm to solve the flow shop problem with incorporated sequence-dependent transportation times with a single-transporter. Several heuristic algorithms were developed considering the transportation of an interrupted job from a machine to another machine. Naderi et al. [18] consider the transportation times in permutation flow shops and provide solution methods including heuristics and metaheuristics to solve large-sized problems. An adaptation of the imperialist competitive algorithm hybridized by a simulated annealing based local search was proposed by Karimi et al. [11] to solve the problem.
FJSP. Schworm et al. [21] investigate a QA-based FJSP on large problem instances specific to modern manufacturing and demonstrate the efficiency of the approach regarding scalability, solutions’ quality, and computing time. Denkena et al. [9] compare the computing time of several approaches to solve the classical FJSP and conclude that it generally increases significantly with the problem size. Their research suggests that these methods deliver valid results on small problem instances, but are not suitable for problem sizes relevant to practice. Therefore, methods are required that yield high quality outcomes, while they also ensure consistently quick computation times, particularly for large problem instances. The paper additionally investigates the concept of QA based optimization for a process parallel FJSP on existing benchmarks. The results reveal that the QA technique consistently achieves low anneal times, especially for large instances, while it still produces outcomes comparable to those of classical heuristics.
DFJSP. Only few papers focus on the FJSP [28] with transportation time between jobs, also known as DFJSP. Zhang et al. [28] develop a modified Genetic Algorithm (GA) to solve the FJSP with transportation time. Marzouki et al. [15] use a Chemical Reaction Optimization Metaheuristic to solve the DFJSP. Liu et al. [13] propose a refined GA and apply it with satisfactory results on a real case of a manufacturing company.
Generally, most of the papers focus on studying the JSSP. Although there is some research that includes transportation time in the model, these use classical approaches to solve the problem, like heuristics or GA. In this research, an extended case of DFJSP is studied, where not only jobs, but also job operations can be distributed. For this purpose, a D-Wave Quantum Annealer is used to solve the DFJSP with distributed operations. The availability of a set of machines capable to process the same operation and the geographical distribution of these machines add complexity and pose a significant challenge to solve this combinatorial optimization problem.
4 Modell
In this Section we describe the QUBO model to solve the extended DFJSP as required by the wool textile industry. To this end, the extended problem is first formally described in Subsection 4.1. Based on this, we create the QUBO model in Subsection 4.2, reduce the number of its variables in Subsection 4.3 and deal with the determination of the Lagrange parameters in Subsection 4.4.
4.1 Problem formulation
The DFJSP refers to job scheduling in distributed manufacturing environments. There could be different objectives when solving the DFJSP (e.g., minimizing job completion time, processing costs or energy consumption). This paper focuses on finding the optimal schedule that minimizes the time required to complete the jobs, i.e., the makespan. The formulation used is the one introduced by Venturelli et al. [24] for the JSSP and the extension from Schworm et al. [21] for the FJSP. In this research, the model is extended to include shipping time for distributed operations. It is assumed that there are A jobs J = {j, …, j} to be processed on B machines M = {m, …, m} in C factories F = {f, …, f}. Each job consists of a set of L operations = {o, …, o}. Any operation can be processed on at least one machine out of the given set M. For any job, the order of operations is given by = {r, …, r}, in which r is the last operation of job . Another notations used in this paper are listed in Tbl. 1.
| Notation | Description |
| o | Next operation of any operation . |
| m | Machine of next operation o. |
| t | Time step in a given timeline T = {0, …, t}, where t is the completion time of all jobs. |
| t | Next time step of any time step . |
| t | Starting time of operation o of job i on machine m. |
| p | Processing time of operation o of job i on machine m. |
| d | Shipping time between machines. |
| k | Shipping time between factories. |
| P | Minimum predecessor time of operation o. |
| S | Minimum successor time of operation o. |
Reminder: The production orders are referred to as jobs and the production steps as operations in the context of the DFJSP. Since machines are associated with factories, the distance between machines is derived from the distance between factories. Distance between factories is directly related to the shipping time needed to send the lots in production to another location. In this paper, distance and shipping time is used interchangeably, and refer to the value that is added to the cost function when production lots are shipped between locations. Fig. 2 shows a schema representing the DFJSP.
Different jobs may have a different number of operations, each with its own processing time on a specific machine. This paper focuses on a special case of DFJSP where operations of the same job can be performed on machines in geographically distributed factories, which means that the distance between machines should be considered when generating the job schedule.
In the DFJSP the following constraints are considered:
1.
Precedence of operations - an operation can only start, when the previous operation has finished.
2.
Operation start once - operation processing on machines cannot be interrupted (the reason behind this condition is that each machine must be technically adjusted for every new production lot, which is time-consuming).
3.
No-overlap - each machine can only process one operation at a time.
Other underlying assumptions of the DFJSP are that operations of a job must be performed in the correct order to complete that job. The operations of different jobs are not subject to any precedence constraints and can be processed in parallel. Each factory is equipped with machines of different types, thus leading to different potential processing times of an operation. A job can be scheduled on a machine more than once. All machines are available for use and all jobs can be processed starting with time zero [28].
4.2 QUBO model
Samplers are processes used to solve problems with D-Wave. A BQM sampler samples low energy states in the model defined by the QUBO equation and returns a list of samples to increase energy. The QUBO model is composed of binary variables x, each represent one operation per unit of time on a machine, up to the makespan:
This variable determines whether an operation o of a job i starts on a machine m at time t or not. Time t goes from time 0 to t in integer steps, where t is the sum of the execution times of all operations and the shipping time between machines.
The objective function H(x) is defined as a cost function (Eq. 5), which minimizes the time taken to complete all operations (the makespan). Cost functions are mathematical expressions of optimization problems. For quantum computing, cost functions are quadratic models that have the lowest values (energy) for good solutions to the problems they represent. Each point on a cost function represents a different solution configuration - in this case, each configuration is a particular assignment of the starting times of operations. The goal of the optimization is to minimize the cost of the solution, i.e. the amount of time needed to complete all operations. The mathematical representation of the penalty terms and of the makespan objective are added and build the cost function of the DFJSP with the format:
where:
f(x), g(x) and h(x) represent penalty functions for precedence, operation once and no-overlap constraints, respectively,
k(x) represents the objective function, which will minimize the makespan,
, , are the Lagrange parameters that emphasize different weights assigned to penalties.
Every time a solution violates one or more constraints, a penalty is applied to that solution (e.g., two operations on a single machine that are scheduled to run at the same time violate the no-overlap constraint, so a penalty will be assigned). Positive terms penalize certain solution configurations, while negative ones support them. By adding penalties to terms that break the constraints, the cost of those configurations increases. This reduces the likelihood that a suboptimal solution will be generated. The importance attributed to each term can be adjusted using the weights , and . The process of adjusting these weights is explained in Section 4.4. The weights determine the importance of penalty functions in relation to other terms.
Precedence constraint. The precedence constraint states that a job operation can only start after the previous operation is finished and the lot arrived at the production site of the next operation if shipping is necessary.
To formulate the penalty, cases where consecutive operations o and o in a job take place out of order are counted. If the starting time of o denoted by t is less than the sum of the starting time of o denoted by t, its processing time p and the shipping time between machines d, then this count as penalty. The penalty condition can be formulated as follows:
This penalty is the sum of all operations of a job i for all the jobs:
where:
sum over all jobs,
sum over all combinations of sequential operations,
penalty condition,
represent the expression that assigns a penalty every time the precedence constraint is violated.
Operation-once constraint. To represent this mathematically, the following function is formulated, which sums values across all possible time steps t:
where:
sum over all jobs,
sum over all operations,
sum over all time steps on all machines,
represent the expression that assigns a penalty every time the precedence constraint is violated,
squares the results in parentheses to keep the penalty term positive in case an operation is not scheduled.
No-overlap constraint. The penalty function is the sum of pairwise products of binary variables and that represent starting time of operations on each machine, which must be equal to zero for the configuration to be valid:
The sets and are defined as follows:
The set A is a condition that forbids operation o to start at time t if the processing of the previous operation o has not finished yet (meaning that ). The set B forbids the operations to start at the same time. Further details on the penalty function are:
sum of each independent machine,
sum over jobs,
sum over operations,
product of binary variables, that assigns a penalty every time the no-overlap constraint is violated.
Makespan Function. To formulate the optimization objective of completing each job in the shortest time possible, the following polynomial function is defined, that penalizes the late completion time of operations:
where:
P is the minimum predecessor time of the operation o, which is the sum of the minimum processing times of the preceding operations and shipping distances between them.
4.3 Variable pruning
In the previous section, the binary variable x was defined to represent whether an operation o of a job i is scheduled to start on a machine m at time t or not. To reduce the number of variables in the model, a method called variable pruning is used, which removes variables that are not possible in valid solutions. First, an upper bound t of the highest possible makespan is calculated. Without this step, the solver will have to consider all solutions between zero to infinity, which implies a very large solution space that makes it difficult to embed and solve the QUBO [24]. Instead, the search space is limited by defining the upper bound of a schedule as the sum of the maximum processing time of each operation in all jobs and maximum shipping time between these operations. Thus, a significant number of variables x can be pruned by applying restrictions on time t.
Next, the duration and order of operations within a job are examined. Variables that allow an operation to execute at a given time that results in an invalid schedule are removed. Specifically, variables are pruned to allow operations to start earlier than the sum of the processing time of its predecessors, or if they allow operations to start later than the sum of its own processing time and that of the successor operations subtracted from t [29].
To set a lower bound of operations in a job, the notion of minimum predecessor time P of an operation is used:
P is the minimum predecessor time of the operation o, calculated as the sum of the minimum processing time of the preceding operations and shipping distance between them. A starting time earlier than the minimum predecessor time is invalid, therefore, these binary variables can be eliminated from the model [21].
Additionally, an upper bound for the starting time t of operations can be set. This corresponds to the maximum makespan t minus the sum of the processing time of the operation and that of all subsequent operations and shipping time between machines S:
where:
These variables can be also eliminated from the QUBO, since they represent invalid starting times [24]. Thus, the starting time of operations o should be in the range , all the other variables that are not in this range are removed from the model to decrease the size of the computational problem and computation time [29].
To further reduce the number of binary variables, the starting times of simultaneous operations on machines are also removed [9].
4.4 Lagrange parameters
In the previous sections the constraints have been formulated as binary polynomials in a Hamilton formulation. The binary polynomials f(x), g(x), h(x) are multiplied with non-negative scalar weights also known as Lagrange parameters (Eq. 5), which determine the impact of the respective polynomial [21]. Weights are assigned to each constraint depending on the relative strengths of the objective function and the constraints. The weights have to be set high enough, so that valid solution with no broken constraints will be returned. On the other hand, values that are too high will result in solutions that are valid, but far from optimal777D-Wave Problem Formulation Guide Whitepaper: https://www.dwavesys.com/media/bu0lh5ee/problem-formulation-guide-2022-01-10.pdf.
To find the Lagrange parameter for the operation-once constraint, the case of H(x) is investigated that consists of the constraint g(x) and the makespan objective function:
For example, for job j in Tbl. 2, operation o is starting on machine m at time t and has a processing time p (this technique is applied in the same way for larger problem instances with more jobs and operations).
| t | |||
| 0 | 1 | 0 | 0 |
| 1 | 0 | 1 | 0 |
| 2 | 0 | 0 | 0 |
| 3 | 0 | 0 | 0 |
| - | 3 | 2 |
Thus, the sum of processing times of previous operations of o is equal to 1. Further, o of the same job is starting at different times t on machine m. Shipping time between machines d equals to zero. Both binary variables x and x are equal to 1. This can be expressed by a QUBO. For , the value and thus, the energy of cost function H(x) equals to 3:
This is the sum of the processing times of the two operations o and o that represent the makespan objective function (second term in equation H(x)). Since the start-once constraint is not violated, the first term in equation H(x) is equal to zero. This is also the minimum energy level of a valid solution for this specific problem instance. Any violation of the constraints will add penalties to the cost function that will increase the energy of the system.
A specific case should, however, be considered. Let us assume that operation o is not scheduled to start at all. This will result in a cost function H(x) that equals to 2:
Although the returned solution is invalid, the energy of the system equals in this case, which is less than the minimum valid energy of . Tbl. 2 summarizes the results for function H(x) for .
To compensate for this effect, a Lagrange parameter is applied that increases the weight of the penalty term. In DFJSP valid solutions with higher energy are preferred to invalid solutions with low energy but violated constraints. The value of the weight parameter depends on the problem instance. A value that is too low penalizes invalid configurations of variables insufficiently, resulting energy values that might be lower than the minimum energy of the best valid solution. Values that are too high would return valid configurations – possible with higher energy than the optimum. The value of the Lagrange parameter is set based on the maximum makespan of the jobs ( t in Section 4.3). The maximum energy value, which corresponds to the maximum makespan t, is equal to 10 (this value is calculated for each problem instance based on the input job dictionary). The penalty is set to , meaning the maximum energy possible for a valid configuration for this problem instance, which results in the following function H(x):
The energy of the function is equal to , which is above the maximum energy possible for valid solutions of . This ensures that no invalid solution will have lower energy than any of the valid solutions. The technique demonstrated in this simplified example can be applied in the same way to problems that have more jobs and operations. The resulting QUBO will have more variables.
The same method is used to calculate the weights for other constraints. Generally, the Lagrange parameters are set individually for each constraint, depending on the magnitude of the objective function to be minimized and its relative strength to the constraints. In this case, to make sure that the cost function H(x) has the lowest energy for valid solutions, the weights of penalty functions f(x), g(x) and f(x) are all set equal to t, which is the maximum energy value of all valid solutions.
5 Experiments
In this Section, we implement the model and conduct experiments on a quantum annealer of D-Wave (more precisely, the AdvantageSystem4.1). As it is common practice, the results of the experiment on a quantum annealer are compared with results obtained from SA. Firstly, the problem size (see 5.1) and solver parameters (see 5.2) are defined. The QUBO model is constructed with PyQUBO Python library, which is a high-level interface that helps formulating combinatorial optimization problems in QUBO form, and supports solving them using various solvers, including D-Wave quantum annealer [27]. We then present the results of the experiments in Subchapter 5.3 and discuss them in Subchapter 5.4.
5.1 Problem size and dataset
To analyze the performance of the D-Wave quantum annealer for the DFJSP, 6 problem instances are defined in such a way that the size of the smallest problem instance is 50 and the size of the largest one is 250. The instances in between are defined in even intervals of 50 (e.g. 100, 150, etc888Problem size is rounded for some instances (e.g. instance size 99 is rounded to 100).). The size of the problem instances refers to the number of binary variables x necessary to represent the problem in QUBO form. The maximum size of the problem instances is dictated by the physical properties of the D-Wave QPU. The D-Wave AdvantageSystem4.1 has 5627 qubits. However, not all qubits are interconnected. When embedding the model variables onto the physical hardware graph of the QPU, in some cases, to represent a single logical variable set of physical qubits called chains are necessary, with couplings between physical qubits capable to realize the correct interaction between logical variables [26]. This leads to embedding patterns that need more physical qubits than there are variables in the original problem.
The purpose of the experiment is to test the capacity and performance of the D-Wave quantum annealer for the DFJSP up to the largest problem that can be embedded into the QPU. Tbl. 3 summarizes the instances solved by QA, the number of qubits needed to embed the model onto the physical QPU and the length of the longest chain in the embedding.
| Problem size | Number of qubits | Max chain |
Smaller problem instances require a lower number of qubits for embedding. It is also true that the length of the chains increases with problem size, which means that with increasing number of variables, it is more difficult to find embedding schemas. Therefore, longer chains are required to connect the variables through couplers. In this model, the largest problem instance that could be embedded was of size 250, that required more than 4000 qubits for embedding. For 300 variables, no embedding could be found.
The problem instances are formulated for 2 jobs, j and j, each job containing a minimum of 2 operations. Each operation can be processed on one machine or a set of two machines, with different processing times on each machine. To include shipping distance in the model, the sets of machines available for each operation include machines from different factories. The consideration of the shipping distance is essential for the processing of Contribution 1. The number of operations in each job, the available machines and the processing time on each machine determine the number of binary variables x necessary to represent the problem in QUBO form.
5.2 Solver parameters
There are 20 solver parameters [6] that can be adjusted when submitting problems to solvers. In this experiment, the default settings are kept for most of them, including the default annealing time of 20s. The parameters that are adjusted in the experiment are the number of reads and the chain strength.
Number of reads. In theory, very long annealing times should keep the system in a low energy state, in practice noise due to hardware imperfections will push the state to a higher energy [4]. Therefore, short annealing times are preferred in order to take many samples from the energy distribution of the problem. The appropriate number of samples depends on the problem itself and on its energy distribution and is usually between 10 and 10000 samples [4]. In this experiment, the number of reads (’num_reads’) is set to 1000. Generally, increasing the number of reads can increase the probability to get better solutions, but improvements diminish due to some problem-dependent numbers [6]. For this model, increasing the number of reads much above 1000 does not seem to be of any advantage. Sampling the problem instance with 200 variables 3000 times produced a solution with an energy of 34 and 2 violated constraints, which is worse than the energy of the best solution of 24 and 1 violated constraint obtained with 1000 reads. This is consistent with the results obtained by Carugno et al. [4], who found that increasing the number of reads from 1000 to 10000 did not have any beneficial effects.
Chain strength. For a solution to be consistent, the chain of qubits has to act as a single qubit, meaning that all the qubits in the chain must have the same values after annealing [4]. Broken chains show different values of the qubits at the end of the annealing process. Chains must be set strong enough to ensure that the qubits of the chain are showing the same value compared to other qubits. If the chains strength is too small, then the qubits will show different values. If the chain strength parameter is too high, this will impede the qubits in the chain to change their state, the chain begins to behave like a separate entity that does not interact with other terms of the function.
We follow the D-Wave recommendation to determine the chain strength for each problem instance and start with two utility functions to estimate the chain strength [7]:
1.
scaled
2.
uniform_torque_compensation
The first function is returning a chain strength that is scaled to the problem bias range. The second function calculates a chain strength that attempts to compensate for the torque that would break the chain. Both methods are to be seen as a starting point to estimate the optimal chain strength, and not as rule, since this parameter is problem-dependent. In the next step, the chain strength parameter found in step 1 should be doubled or tripled, to check if this change has an impact on the percentage of broken chains.
To find the best chain strength for the model, the model is run on the QPU for each problem instance, and the resulting energy and the percentage of broken chains in the best solution are analyzed. The intention is to find the lowest energy possible, to at the same time having a low percentage of broken chains to ensure a consistent solution. Fig. 3 shows the results of the measurements for the problem instance of size 150.
The utility function scaled returns a chain strength of 32 with an energy of 69, the function uniform_torque returns a value of 309 with an energy of 966. To understand the relation between these parameters, the range between these two chain strengths is divided into 11 intervals and the energy and the percentage of broken chains in the best solutions for each of these points is measured. Additionally, the smallest chain strength value is divided by 2, to check whether a smaller parameter could improve the results. The results in the graph show that the lowest energy achieved corresponds to the next point in the interval from the point obtained with the scaled function. However, the percentage of broken chains corresponding to this value is 41%, which is high compared to values in the graph for higher chain strength values. In order to decide whether to accept a chain strength parameter of 39 or not, an additional analysis of the consistency of the solution is necessary. Fig. 4 shows that, despite a higher number of broken chains, the embedding with a chain strength of 39 delivers a consistent solution without broken constraints. With the increasing chain strength parameter the number of broken constraints increases too.
This procedure is repeated in the experiment for all problem instances to calculate the appropriate parameter for the chain strength. For all instances, the parameter that delivered the best solution with the lowest energy is one interval above the parameter estimated by the scaled function. Since the intention is to test to which extent the QPU is able to handle large problems, this method is applied to problem instances of size 250 and 300. For problems of size 50 to 150, the solver is able to find consistent solutions with no broken constraints. For problem instances of size 200 and 250, the solutions with the lowest energy found had 1 and 2 broken constraints, respectively. For the instance of size 300 no embedding could be found.
SA parameters. To solve problems with SA, the D-Wave Sampler class SimulatedAnnealingSampler of the D-Wave Ocean SDK is used. The sampler utilizes a classical heuristic SA algorithm that runs locally on the CPU999Hardware configuration: 2x Intel Xeon Gold 6326 CPU 16Core, 2.9GHz 24M 185W Socket 4189, 4 x Samsung 64GB DDR4 ECC RDIMM 3200Mhz total 256GB RAM, 4 x HDD WD Gold Datacenter 1TB SATA, 1x SSD Samsung PM883 480GB SATA, 2 x 1 / 10 GB Ethernet 10 BaseT RJ-45 onboard.. For SA, the Central Processing Unit (CPU) time is measured, which is the amount of time that the CPU has spent solving the problem. The CPU time is compared to QPU access time.
To compare QA and SA, the number of iterations needed and the ability to find a consistent solution are analyzed. To have comparable results, the number of reads for SA is set to 1000, like the parameter for QA. The parameter number of sweeps (num_sweeps) is set to 3, which is the smallest value for this parameter that returned solutions with no broken constraints. As with QA, the SA algorithm is run for all the instances of the problem to determine whether the system is able to find valid solutions. Additionally, problem instances of size 350 and 400 are included in the analysis.
5.3 Numerical Results
In order to provide the Contributions 2 and 3 formulated in Section 1, the results obtained from the experiment are analyzed from both a qualitative and a quantitative perspective. The first answers the question, how good the returned solution is in terms of system energy and makespan. The second compares QPU access time and CPU time of SA to assess whether there is a speed advantage in using the quantum annealer to solve the DFJSP with distributed operations. The results are influenced by the hardware characteristics, on different hardware other results might be obtained.
System energy and makespan. The goal of the experiment is to find consistent solutions in the shortest time possible. When sampling from QA or SA, the returned sample set contains 1000 samples, in accordance with the defined number of reads. The sample with the lowest energy in the sample set is selected as the best sample result for the respective problem instance. This is not always the solution corresponding to the global minimum represented by the lowest energy possible for the QUBO. For this use case, however, finding the global minimum is not crucial. A solution is satisfactory if it is consistent and represents a valid schedule plan obtained in a reasonable time. Further optimization steps can then be implemented to minimize the energy of the system and the makespan. Fig. 5 shows a comparison of best solution energies found with QA and SA for instances of different sizes.
For the smallest problem instance, both QA and SA found a low energy of 2 and returned valid solutions. For instances of size 100 and 150, both methods delivered consistent solutions, but SA performed better in terms of energy. For problem instances of size 200 and 250, the best solution returned by quantum annealer had 1 and 2 broken constraints, respectively. This explains the increasing energy values for these problems. This observation is related to the percentage of broken chains that increase with problem size and deteriorate solution quality with the growing number of variables in the problem. These results are consistent with results found in other papers. Carugno et al. [4] found that for JSSP instances with an unknown optimal solution, QA is not able to find the best solution that corresponds to the global minimum.
To calculate the makespan in the makespan function (Eq. 10), the minimum ending time and the maximum ending time of all operations are used. These terms represent the minimum and the maximum makespans possible and are the lower and the upper bound of the makespan range, respectively. In terms of makespan, QA returns solutions in the lower half of the possible makespan range for all problem sizes, except for size 250 that is in the upper half of possible solutions (Fig. 6).
To summarize, QA is able to find consistent solutions for problem instances up to 150 variables. For the instance of size 200, the makespan found is still in the first half of solutions, but 1 constraint is violated. For size 250, the makespan is slightly below the maximum possible value and 2 constraints are violated. Current QA technology is subject to hardware and software limitations, but with the growing number of qubits and increasing connectivity between them, problems with a higher number of variables can be potentially embedded into future QPUs. Additionally, improving QPU stability promises to deliver solutions of better quality and with lower energy of the system.
Next, QPU access time and CPU time of SA for the DFJSP with distributed operations are compared.
Calculation time. To determine the QPU access time, the instances from size 50 to 250 are solved with the quantum annealer. Based on these results, a trendline for instances of size 300 up to 400 is derived. In this case, since the rate of change in the data decreases and almost levels out, a logarithmic trendline describes the data the best. The R-square value of 0.9767 is a strong indicator that the logarithmic trendline corresponds to the data. With SA it was possible to solve problem instances of all sizes up to the problem instance of 400 variables. Fig. 7 shows the results that were obtained by both methods.
Although SA performs better for small problem sizes, the CPU calculation time of SA increases exponentially with the increasing problem size. On the other hand, due to hardware limitations, the AdvantageSystem4.1 quantum annealer is capable to embed problems up to a maximum size of 250 variables. However, its advantage is that the calculation time increases at a decreasing rate with problem size. For business problems with a very large number of variables, specific to the DFJSP, a calculation time that changes logarithmically with problem size is a big benefit, compared to the exponential growth of the CPU time.
5.4 Discussion
In this paper, a step-by-step practical implementation of the DFJSP with distributed operations and its evaluation using D-Wave quantum annealer, based on the real use case from the wool textile industry is presented. First, the problem is described mathematically as a constraint satisfaction problem. Then, the process of reformulating the mathematical problem in QUBO format is discussed in detail, as well as its embedding into the D-Wave quantum annealer QPU and the configuration of QPU solver parameters. Finally, solutions are sampled from the QPU to find low-energy states of the cost function. While formulating the QUBO, particular attention is dedicated to the calculation of the appropriate Lagrange parameters (Subsection 4.4), that increase the weights of the penalty terms. Another aspect that is carefully investigated while configuring the experiment is the chain strength parameter (Subsection 5.2). Both aspects make up Contribution 2.
Another aim of the paper was to investigate whether QA is able to provide a speed advantage for the DFJSP with distributed operations specific to the wool textile industry. To answer this question, both qualitative and quantitative analyses are conducted. Current QA technology allows more qubit usage than classical gate model quantum computers. However, there are still limitations imposed by hard- and software imperfections. With this in mind, the test challenged up to the limit the ability of QA to deliver consistent solutions for large problem instances in the shortest time possible. To conduct the experiment, problem instances from size 50 to 250 are solved in steps of 50 variables with QA and with SA. Additionally, problems up to 400 variables are solved only with SA to increase the size range and determine the calculation time trend. The missing data for QA is extrapolated as a logarithmic trendline.
The qualitative analysis of the solutions returned indicates that for problem instances up to a size of 150 variables, both QA and SA, with technical characteristics as used in this research, deliver consistent solutions, and thus valid production schedules. However SA performs slightly better in terms of returned energy. With an increasing number of variables, hardware imperfections become visible, which manifest themselves in an increased number of broken chains in the returned solutions, and consequently, in solutions with higher energy due to broken constraints. For the problem of size 200, the best solution returned by QA has 1 broken constraint, for the instance of 250 variables 2 constraints were violated. Due to a higher usage of physical qubits needed compared to the number of variables in the original QUBO, the maximum problem instance that could be embedded onto the QPU was of 250 variables. For the problem of size 300 no embedding could be found. In terms of makespan, QA finds solutions for almost problem sizes in the lower half of the possible makespan range, except for size 250 which is the upper half of possible solutions. The makespan determined by SA is comparable, although for a few problems instances SA is able to find solutions that are closer to the global minimum. All solutions calculated by SA are consistent, even for the largest instances in the test.
To tackle the Contribution 3, QPU access time and SA CPU time necessary to calculate the solutions to the problems were measured. The results reveal that QA needs more time than SA for smaller problem instances, but the time needed increases at a decreasing rate with the growing problem size. This is not the case for SA, whose CPU time grows exponentially with the increasing number of variables. The CPU timeline and the trendline of the CPU access time are slightly above the value of 300 variables. If this trend will persist with the new QA hardware generation that is to come, every problem with a size larger than this value would potentially be solved faster than with SA. Thus, the question of the speed advantage can be answered positively: QA can potentially result in a speed advantage for the DFJSP with distributed operations.
Even if the QA technology currently still faces limitations, the new generation of quantum annealers with improved hardware promises to speed up solving complex combinatorial optimization problems like the DFJSP. The measurements show that, given the same experimental parameters and improved hardware, QA could provide a speed-up advantage over SA starting with problem instances slightly larger than 300 variables. To test this assumption, new tests must be performed for larger problem instances, for which only the trend could be extrapolated at the moment, as soon as the new generation of quantum annealers becomes available. More qubits and connections between them, as well as more stable systems, will allow embedding larger problems and potentially finding energies closer to the global minimum. Additionally, research on tuning and configuration of QA parameters is another necessity that will allow to minimize the energy and find the best solution. Determining the strength of the qubit chains is a laborious task that is not only model-dependent, but also instance-dependent, and has a large impact on the quality of the obtained solutions.
6 Conclusion
Many modern manufacturing enterprises evolved from a single to a multi-factory production environment, as is the case with big production companies in the wool textile industry. Planning systems changed from classical JSSP to DFJSP, with companies facing the challenge of increased frequency and complexity of production planning. Classical heuristics deliver valid results on small problem instances but are unable to find solutions for problem sizes relevant to practice. Therefore, methods are required that yield high quality outcomes while reasonable computation times are ensured. In this paper, the concept of QA-based optimization for the DFJSP with geographically distributed operations is investigated to answer the question whether QA is able to provide a speed advantage for the DFJSP with distributed operations specific to the wool textile industry.
To do this, we show a step-by-step implementation of the workflow for solving the DFJSP with QA (Contribution 1). A special attention is devoted to topics that do not have a straightforward approach but a great impact on the solution quality, such as calculating the Lagrange parameters or determining the chain strength. The results show that up to a certain problem size, QA delivers consistent solutions that are not the global minimum solution in the solution space but yield a makespan in the first half of the allowed valid solutions (Contribution 2). For all instances, SA delivers comparable or slightly better results, which are consistent with other research. In terms of calculation time, QA needs more QPU time to find solutions for smaller problem instances in the test, but QPU access time increases at a declining rate with the growing problem size. In contrast, the SA CPU time develops exponentially for problems with a larger number of variables. The results of the experiment show that starting with a certain problem size, QA could potentially offer a speed-up advantage, given that the expected improvements in hard- and software will take place (Contribution 3).
One of the future tasks is to solve larger problems on the next generation of quantum annealers to confirm the results obtained in this experiment. Other topics open for future research are QPU parameter tuning and configuration. There is also a need for additional research on embedding techniques that could minimize the length of the chains and allow the embedding of more QUBO variables into physical qubits, thus solving problems of greater size.